OpenStudy (anonymous):

Anna went to the grocery store. She bought 3 oranges and 6 apples. She paid $4.50. Maria went to the same store and bought 2 oranges and 2 apples. Her cost was $2.10. How much would one apple cost?

OpenStudy (ivettef365):

x = oranges y = apples so you have the following equations 3x + 6y = 4.50 2x + 2y = 2.10

OpenStudy (ivettef365):

can you solve the system of equation

OpenStudy (anonymous):

no

OpenStudy (ivettef365):

now you can add as follows: 3x + 6y = 4.50 -6x - 6y =-6.30 ============ -3x = -1.80

OpenStudy (ivettef365):

now divide by -3 and you have x = .60 so the oranges cost 0.60, find how much the apples cost

OpenStudy (ivettef365):

do you know how to ?

OpenStudy (ivettef365):

substitute the x = .60 on the equation so you have 3 (.60) + 6y = 4.50

OpenStudy (ivettef365):

can you solve for y